OpenStudy (anonymous):

A person want to invest a certain amount at 65% p.a. S.I. If he wishes to receive a monthly interest of 3000, what is the amount to be invested?

OpenStudy (ribhu):

si=prt/100

OpenStudy (ribhu):

3000*12 = p*65*1/100 p can be calculated

OpenStudy (anonymous):

Answer options are a:6,00,000, b: 50,000, c: 4,00,000, d: 3,00,000.